# Low-Acceleration Gravitational Anomaly from Bayesian 3D Modeling of Wide Binary Orbits: Methodology and Results with Gaia DR3

### v3

@@ Abstract
Isolated wide binary stars provide natural laboratories to directly probe gravity for Newtonian acceleration $g_{\rm{N}}\lesssim 10^{-9}$ m s-2. Recent statistical analyses of wide binaries have been performed only with sky-projected relative velocities $v_{p}$ in the pairs. A new method of Bayesian orbit modeling exploiting three relative velocity components including the radial (line-of-sight) component $v_{r}$ is developed to measure a gravitational anomaly parameter $\Gamma\equiv\log_{10}\sqrt{\gamma_{g}}\equiv\log_{10}\sqrt{G_{\rm{eff}}/G_{\rm{N}}}$ where $G_{\rm{eff}}$ is the effective gravitational constant for pseudo-Newtonian elliptical orbits, while $G_{\rm{N}}$ is Newton’s constant. The method infers individual probability distributions of $\Gamma$ and then combines the independent distributions to obtain a consolidated distribution in a specific range of $g_{\rm{N}}$. Here the method is described and applied to a sample of 312 wide binaries in a broad dynamic range $10^{-11.0}\lesssim g_{\rm{N}}\lesssim 10^{-6.7}$ m s-2 with $v_{r}$ uncertainties in the range $380168<\sigma_{v_{r}}<$ 380 m s-1 selected from the Gaia DR3 database. The following results are obtained: $\Gamma=0.000\pm 0.011$ or $\gamma_{g}=1.00\pm 0.05$ ($N_{\rm{binary}}=125$) for a high acceleration regime ($10^{-7.9}\lesssim g_{\rm{N}}\lesssim 10^{-6.7}$ m s-2) agreeing well with Newton, but $\Gamma=0.085\pm 0.040$ or $\gamma_{g}=1.48_{-0.23}^{+0.33}$ (35) for a MOND regime ($10^{-11.0}\lesssim g_{\rm{N}}\lesssim 10^{-9.5}$ m s-2) and $\Gamma=0.063\pm 0.015$ or $\gamma_{g}=1.34_{-0.08}^{+0.10}$ (111) for a MOND+transition regime ($10^{-11.0}\lesssim g_{\rm{N}}\lesssim 10^{-8.5}$ m s-2). These results show that gravitational anomaly is evident for $g_{\rm{N}}\lesssim 10^{-9}$ m s-2 and $\Gamma$ in the MOND regime ($\lesssim 10^{-9.5}$ m s-2) agrees with the first-tier prediction ($\Gamma\approx 0.07$ or $\gamma_{g}\approx 1.4$) of MOND gravity theories.
@@ 1 Introduction
Several statistical analyses of wide binary internal motions carried out during 2023 - 2024 based on Gaia data release 3 (DR3; Vallenari et al., 2023), by Chae (Chae, 2023a, 2024a, 2024c) and independently by Hernandez et al. (Hernandez, 2023; Hernandez et al., 2024a; Hernandez & Kroupa, 2025), indicate that standard gravity is not only broken at low internal Newtonian acceleration ($g_{\rm{N}}\lesssim 10^{-9}$ m s-2) but gravity is boosted by approximately 40% at $g_{\rm{N}}\lesssim 10^{-10}$ m s-2 in agreement with the generic prediction of modified gravity theories (Bekenstein & Milgrom, 1984; Milgrom, 2010) in the paradigm of modified Newtonian dynamics (MOND; Milgrom, 1983). Although divergent results (Banik et al., 2024) were published during the same period, the recent critical review by Hernandez et al. (2024b) (see also Chae 2024c for more details) addresses the origin of the divergent results, in particular the lack of the self-calibration procedure for the fraction of hierarchical systems in an impure binary sample (i.e. including hierarchical systems).
@@
All recently published results are based on the relative sky-projected (i.e. transverse) velocities $v_{p}$ in the pairs rather than the relative physical velocities $\mathbf{v}$ (or its magnitude $v$). This is because the Gaia DR3 database provides precise proper motions and parallaxes (so precise $v_{p}$) but far less precise or no relative radial (i.e. line-of-sight) velocities $v_{r}$. Existing tests with $v_{p}$ alone are less direct than using $v$ itself, and thus involve more complications. Furthermore, gravity cannot be individually constrained for each binary, but only a statistical population can be compared with the corresponding population predicted by a gravity model in conjunction with statistical correlations/assumptions of the binary population (see Chae 2023a).
@@
For an ideal individual measurement of gravity (with known mass or vice versa) with a binary system, its orbit would have to be traced for a significant fraction of the period with well-measured $\mathbf{v}$ values. In reality, due to the long periods $\sim 10^{5-6}$ yr for wide binaries of interest (see, e.g., Figure 2 of Chae 2024a), we will have to work essentially with one value of $v$ at a random phase of the orbit. It is then unclear how well gravity can be individually constrained by such a snapshot observation of the orbital motion even if $\mathbf{v}$ is well measured. One may imagine that any individual inference of gravity will cover a broad range of possibilities, if any meaningful inference is feasible at all. To address this question in search of a more direct test of weak gravity than existing analyses, here I develop a Bayesian orbit modeling method with snapshot measurements of wide binary internal motions for hundreds of exceptional systems from the Gaia DR3 as well as mock wide binaries with controlled precision of $v_{r}$ inspired by ongoing and future observations. Note that, as details will be given below, even the exceptional systems from the Gaia DR3 have much less precise $v_{r}$ than $v_{p}$.
@@
Not surprisingly, as will be shown in this work, Bayesian inferred individual probability density functions (PDFs) of gravity with the Gaia DR3 database are too broad to distinguish gravity theories of current interest (in particular, Newton and MOND). This will, in general, be true even for future data with much more precise $v_{r}$ due to the snapshot nature of the observations. However, when individual PDFs represent independent inferences of the same quantity (a gravity parameter in the present case) with different data, the PDFs may be combined to maximize the information from all data (Hill, 2011; Hill & Miller, 2011). With such a statistical combination/consolidation method called “conflation”, we will see that even the current data clearly indicate a gravitational anomaly in the low-acceleration regime, in line with the results of 2023-2024 based on $v_{p}$ published by Chae and Hernandez et al. This holds great promise for future studies with a significant number of precise values of $v_{r}$.
@@
This paper is structured as follows. In Section 2, I describe the methodology and data. In detail, I first develop a Bayesian methodology to infer probability density functions (PDFs) of gravity from individual binaries (Section 2.1), and then describe the method of statistically combining individual PDFs to obtain a consolidated probability distribution by maximizing information (Section 2.2). Section 2.3 describes the selection of a pure binary sample from the Gaia DR3 database that will be used for individual Bayesian modeling. The selection of the sample will be more stringent than previous samples that were used mainly for median statistics in bins of accelerations (or equivalent variables) by Chae (2024a, c). In Section 3, I describe the results of Bayesian modeling based on the Gaia sample. In Section 3.1, I first examine the individual Bayesian results and the overall properties of the results. In Section 3.2, I present the consolidated probability distributions of gravity with wide binaries in a Newtonian (relatively strong internal acceleration) regime. In Section 3.3, I present the main results of this work, i.e. the consolidated probability distributions of gravity in low-acceleration regimes. In Section 4, I discuss the meaning of the results and future wide binary gravity research. In Section 5, I present the current conclusion and prospects with future data on radial velocities. In Appendix A, a general transformation with the Euler angles between the orbital plane and the observer’s frame is described as it is needed to produce mock samples. In Appendix B, perspective effects in proper motions and radial velocities of wide binary stars are investigated to explore any possible biases in the low-acceleration regime data.
@@ 2.1 Bayesian 3D modeling
Gravitational anomaly with respect to the prediction of Newtonian gravity can be defined and tested/measured in various ways. Because of the extremely long periods of wide binaries compared with the observation time span, only one relative velocity is observed in effect. In Newtonian gravitational dynamics, the relative motion between the pair in an isolated binary system follows an ellipse. In nonstandard theories of gravity such as MOND gravity (e.g., Bekenstein & Milgrom 1984; Milgrom 2010), the orbits of wide binaries under the external field of the Milky Way oriented at an arbitrary angle with respect to the orbital plane may not be exact ellipses. However, whether actual orbits are exact ellipses or not, I will work with elliptical orbits. The assumption of elliptical orbits will have no impact on testing Newtonian gravity, i.e. whether a gravitational anomaly is present or not. However, when a gravitational anomaly is present, the quantified gravitational anomaly must be understood as an effective quantity for elliptical orbits.
@@
Consider Figure 1 that connects the observer’s primed frame with the unknown orbital plane. Note that the $x$ and $x^{\prime}$ axes coincide by the definition of the two coordinate systems in this figure. (See Appendix A for fully general transformations.) As in Gaia observations, suppose that observed quantities for the two stars of each binary are the R.A. ($\alpha_{k}$ with $k=A,B$ referring to the brighter and fainter components hereafter) and Decl. ($\delta_{k}$) given in units of degree, PMs ($\mu^{\star}_{\alpha,k}$ and $\mu_{\delta,k}$) given in units of mas yr-1, and radial velocities ($\text{RV}_{k}$) given in units of km s-1. What matter for internal dynamics are the relative positions and relative velocities between the pair.
@@
which will be used as an observational constraint on the orbit. The relative radial velocity is given by
@@
The above equations (Equations (5) and (7)) for the relative velocity components based on the observed PMs and RVs do not include corrections for the “perspective effects” (Shaya & Olling, 2011) that are apparent relative velocities between the pair in the observer’s spherical coordinates due to the motion of the barycenter of the system in the 3D space. For wide binaries in the solar neighborhood, perspective effects can be relevant for $s>0.1$ pc (i.e. 20 kau) (El-Badry, 2019).
Almost all binaries considered in this work have $s<20$ kau (99% satisfy $s<15.5$ kau) meaning that perspective effects are minimal, and thus their corrections are not warranted when RV uncertainties are large. Moreover, for an accurate estimate of the perspective effects, the difference in distances of the two stars is needed. Unfortunately, observed distance uncertainties are (usually much) larger than $s$ and one has to assume zero (or reasonably small) difference in the distances to obtain a first-order estimate of the perspective effects. The default choice of the present work will be to ignore any perspective effects. However, I will also consider the case of correcting for first-order perspective effects estimated by a numerical procedure described in Appendix B.
I define the ratio of the radial velocity to a projected velocity component as
@@
which is related to the inclination angle $i$ by $\tau_{\rm{obs}}=\tan i$. Finally, the magnitude of the relative physical velocity in the 3D space is then given by
@@
In Equations (10), (12), and (13), $e$ is the eccentricity, $i$ is an inclination parameter, [^1] I note that this parameter is not the same as the angle between the orbital plane and the plane of the sky. $\phi_{0}$ is the longitude of the periastron, and $\phi$ is the phase angle given by
@@
where $\{X_{j,\rm{mod}}(\mathbf{\Theta})\}$ ($j=1,2,3$) are the model predictions given by Equations (10), (12), and (13) and $\{X_{j,\rm{obs}}\}$ are the three observational constraints given by Equations (9), (6), and (8). In Equation (15), $\{\sigma_{j}\}$ are the observational uncertainties estimated as follows.
@@
where $v_{x^{\prime}}$ and $v_{y^{\prime}}$ are the relative projected velocity components given by Equation (4) while $\sigma_{v_{x^{\prime}}}$ and $\sigma_{v_{y^{\prime}}}$ are their uncertainties arising from PM uncertainties, and $v_{r}$ is the relative radial velocity given by Equation (7) while $\sigma_{v_{r}}$ is its uncertainty arising from the uncertainties of RVA and RVB. The above uncertainties are based on the standard error propagation procedure.
@@
For eccentricity $e$ I consider the range $0.9990.001<e<$ 0.999 with a prior probability distribution. For the prior, I consider three choices: a flat prior, a “thermal” probability density $p(e)=2e$, or a separation($s$)-dependent probability density $p(e)=(1+\alpha)e^{\alpha}$ with $\alpha=\alpha(s)$ from Hwang et al. (2022) (hereafter referred to as the “H2022” prior) as described by Equation (18) of Chae (2024a). If the data can determine the orbit well, the flat prior may be sufficient. However, with snapshot data, $e$ is ill-constrained in particular for relatively large uncertainties of $v_{r}$ as in many cases based on the current Gaia DR3 database. When $e$ is ill-constrained (or essentially unconstrained in some cases), the prior matters because the flat prior would be equivalent to assuming that the median eccentricity is $0.5$ while the thermal prior has a median of $\sqrt{0.5}(\approx 0.71)$ more consistent with the observed statistical properties. The H2022 prior has $\alpha\approx 1.3$ for $s>1$ kau(kilo astronomical unit) implying a median of $\approx 0.74$ that differs from the thermal case by just $0.03$, indicating that two priors would make little difference. Moreover, because individual data themselves have some constraining power in this work, an eccentricity probability distribution plays a weaker role as a prior than a fixed input as in previous studies. For this reason, the thermal prior will be the default choice.
@@
For $f_{M}$, I assume a Gaussian probability density for $\log_{10}(f_{M})$ with $(\mu,\sigma)=(0,0.021)$ allowing a 5% scatter of $M_{\rm{tot}}$, which covers uncertainties of mass estimates (see Section 2.3). For $i$ and $\Gamma$, I assume flat priors in the ranges: $-90^{\circ}<i<90^{\circ}$ and $-1<\Gamma<1$.
@@
In this subsection, the Bayesian method described in Section 2.1 is applied to mock Newtonian binary systems to investigate the properties of individual PDFs of $\Gamma$ (Equation (1)). Also, I investigate how individual PDFs of $\Gamma$ based on independent data (i.e. snapshot motions) can be combined to give a consolidated probability distribution.
For the orbital motion shown in Figure 1, consider binary systems with total mass $M_{\rm{tot}}=2.0M_{\odot}$, eccentricity $e=0.75$, and $\phi_{0}=-135^{\circ}$. For the semi-major axis of the elliptical orbit, I consider two cases: $a=0.8$ kau and $8$ kau to represent binaries in the Newtonian ($g_{\rm{N}}\gtrsim 10^{-8}$ m s-2) and MOND ($g_{\rm{N}}\lesssim 10^{-9.5}$ m s-2) regimes. Consider a viewing angle $i=60^{\circ}$ in the geometry of Figure 1. For the sake of illustration, assume that both binaries obey Newtonian gravity regardless of their separation (i.e. $\gamma_{g}=1$ or $\Gamma=0$).
@@
$v(r)$ $=$ $\sqrt{\frac{\gamma_{g}GM_{\rm{tot}}}{r}\left(2-\frac{r}{a}\right)}$ (23)
$=$ $941.9\sqrt{\gamma_{g}\frac{M_{\rm{tot}}/M_{\odot}}{r/\text{kau}}\left(2-\frac{r}{a}\right)}\text{ m s}^{-1},$
@@
where
@@
Figure 2: Each row shows a simulated Newtonian orbit with eccentricity $e=0.75$ and semimajor axis $a=0.8$ or 8 kau. All positions and velocities represent relative displacements and relative velocities between the pair of stars with a total mass of $M_{\rm{tot}}=2M_{\odot}$. See Figure 1 for the geometry and the meaning of the other parameters. For each orbit the colored big dots represent 20 mock snapshot observations with uniformly-spaced $t/P$ where $t$ is the time from the passage of the periastron (at $\phi=\phi_{0}$) and $P$ is the period. Because of long periods only snapshot observations are possible. The sky-projected velocities $v_{x^{\prime}}$ and $v_{y^{\prime}}$ would be observed as PMs in real observations while the line-of-sight (radial) velocity $v_{z^{\prime}}$ would be observed by spectroscopic observations.
@@
Figure 3: Each row shows three examples of the posterior probability density functions (PDFs) of the pseudo-Newtonian model parameters including $\Gamma$ (Equation (1)) derived for the mock Newtonian (i.e. with $\Gamma=0$) data shown in Figure 2. For the Bayesian modeling, an uncertainty of 200 m s-1 is assumed for each radial velocity of a star, thus $\approx 283$ m s-1 for $v_{r}$.
@@
Figure 3 shows examples of Bayesian inference with the above procedure for the mock Newtonian data of the two binaries with $a=0.8$ kau or $8$ kau shown in Figure 2. These results represent some expectations for the current nominal Gaia statistical sample if gravity is Newtonian (or not too different from it). In general, the PDFs of the parameters except for $f_{M}$ are quite broad as expected considering the input sizable uncertainties (i.e. $\sigma_{v_{r}}\approx 283$ m s-1) of the relative radial velocity ($v_{r}$). This is particularly true for the data of the binary with $a=8$ kau because its $v_{r}$ values have absolute values even smaller than $\sigma_{v_{r}}$ in some mock observation phases. The PDF of $\log_{10}(f_{M})$ is always equivalent to the prior Gaussian distribution because of its degeneracy with $\Gamma$, which has a flat prior in a broad range.
Although the individual PDFs of $\Gamma$ and the orbital and geometric parameters $e$, $\phi_{0}$, and $i$ are so broad, they are all consistent with the input values within reasonable confidence limits (e.g. $2\sigma$). Each column of Figure 4 shows the individual PDFs of $\Gamma$ for the mock data set of the Newtonian binary with $a=0.8$ kau or $8$ kau. The lower rows use the logarithmic scale for probability density to enhance visibility.

### v4

# Low-Acceleration Gravitational Anomaly from Bayesian 3D Modeling of Wide Binary Orbits: Methodology and Results with Gaia Data Release 3

@@ Abstract
Isolated wide binary stars provide natural laboratories to directly probe gravity for Newtonian acceleration of $g_{\rm{N}}\lesssim 10^{-9}$ m s-2. Recent statistical analyses of wide binaries have been performed only with sky-projected relative velocities $v_{p}$ in the pairs. A new method of Bayesian orbit modeling exploiting three relative velocity components including the radial (line-of-sight) component $v_{r}$ is developed to measure the gravitational anomaly parameter $\Gamma\equiv\log_{10}\sqrt{\gamma_{g}}\equiv\log_{10}\sqrt{G_{\rm{eff}}/G_{\rm{N}}}$ where $G_{\rm{eff}}$ is the effective gravitational constant for pseudo-Newtonian elliptical orbits, while $G_{\rm{N}}$ is Newton’s constant. The method infers individual probability distributions of $\Gamma$ and then combines the independent distributions to obtain a consolidated distribution in a specific range of $g_{\rm{N}}$. Here the method is described and applied to a sample of 312 wide binaries in a broad dynamic range of $10^{-11.0}\lesssim g_{\rm{N}}\lesssim 10^{-6.7}$ m s-2 with $v_{r}$ uncertainties in the range $380168<\sigma_{v_{r}}<$ 380 m s-1 selected from the Gaia Data Release 3 database. The following results are obtained: $\Gamma=0.000\pm 0.011$ or $\gamma_{g}=1.00\pm 0.05$ ($N_{\rm{binary}}=125$) for a high acceleration regime ($10^{-7.9}\lesssim g_{\rm{N}}\lesssim 10^{-6.7}$ m s-2) agreeing well with Newton, but $\Gamma=0.085\pm 0.040$ or $\gamma_{g}=1.48_{-0.23}^{+0.33}$ (35) for the modified Newtonian dynamics (MOND) regime ($10^{-11.0}\lesssim g_{\rm{N}}\lesssim 10^{-9.5}$ m s-2) and $\Gamma=0.063\pm 0.015$ or $\gamma_{g}=1.34_{-0.08}^{+0.10}$ (111) for the MOND+transition regime ($10^{-11.0}\lesssim g_{\rm{N}}\lesssim 10^{-8.5}$ m s-2). These results show that gravitational anomaly is evident for $g_{\rm{N}}\lesssim 10^{-9}$ m s-2 and $\Gamma$ in the MOND regime ($\lesssim 10^{-9.5}$ m s-2) agrees with the first-tier prediction ($\Gamma\approx 0.07$ or $\gamma_{g}\approx 1.4$) of MOND gravity theories.
@@ 1 Introduction
Several statistical analyses of wide binary internal motions were carried out during 2023 - 2024 based on Gaia Data Release 3 (DR3; Vallenari et al., 2023), by Chae (Chae, 2023a, 2024a, 2024c) and independently by Hernandez et al. (Hernandez, 2023; Hernandez et al., 2024a; Hernandez & Kroupa, 2025), wich indicate that standard gravity is not only broken at low internal Newtonian acceleration ($g_{\rm{N}}\lesssim 10^{-9}$ m s-2) but gravity is boosted by approximately 40% at $g_{\rm{N}}\lesssim 10^{-10}$ m s-2 in agreement with the generic prediction of modified gravity theories (Bekenstein & Milgrom, 1984; Milgrom, 2010) in the paradigm of modified Newtonian dynamics (MOND; Milgrom, 1983). Although divergent results were published during the same period, the recent critical review by Hernandez et al. (2024b) (see also Chae 2024c for more details) addresses the origin of the divergent results, in particular the lack of the self-calibration procedure for the fraction of hierarchical systems in an impure binary sample (i.e., including hierarchical systems).
@@
All recently published results are based on the relative sky-projected (i.e., transverse) velocities $v_{p}$ in the pairs rather than the relative physical velocities $\mathbf{v}$ (or its magnitude $v$). This is because the Gaia DR3 database provides precise proper motions and parallaxes (so precise $v_{p}$) but far less precise or no relative radial (i.e., line-of-sight) velocities $v_{r}$. Existing tests with $v_{p}$ alone are less direct than using $\mathbf{v}$ itself, and thus involve more complications. Furthermore, gravity cannot be individually constrained for each binary, but only a statistical population can be compared with the corresponding population predicted by a gravity model in conjunction with statistical correlations/assumptions of the binary population (see Chae 2023a).
@@
For an ideal individual measurement of gravity (with known mass or vice versa) with a binary system, its orbit would have to be traced for a significant fraction of the period with well-measured $\mathbf{v}$ values. In reality, due to the long periods $\sim 10^{5-6}$ yr for wide binaries of interest (see, e.g., Figure 2 of Chae 2024a), we will have to work essentially with one value of $\mathbf{v}$ at a random phase of the orbit. It is then unclear how well gravity can be individually constrained by such a snapshot observation of the orbital motion even if $\mathbf{v}$ is well measured. One may imagine that any individual inference of gravity will cover a broad range of possibilities, if any meaningful inference is feasible at all. To address this question in search of a more direct test of weak gravity than existing analyses, here I develop a Bayesian orbit modeling method with snapshot measurements of wide binary internal motions for hundreds of exceptional systems from the Gaia DR3 as well as mock wide binaries with controlled precision of $v_{r}$ inspired by ongoing and future observations. Note that, as details will be given below, even the exceptional systems from the Gaia DR3 have much less precise $v_{r}$ than $v_{p}$.
@@
Not surprisingly, as will be shown in this work, Bayesian-inferred individual probability density functions (PDFs) of gravity with the Gaia DR3 database are too broad to distinguish gravity theories of current interest (in particular, Newton and MOND). This will, in general, be true even for future data with much more precise $v_{r}$ due to the snapshot nature of the observations. However, when individual PDFs represent independent inferences of the same quantity (a gravity parameter in the present case) with different data, the PDFs may be combined to maximize the information from all data (Hill, 2011; Hill & Miller, 2011). With such a statistical combination/consolidation method called “conflation”, we will see that even the current data clearly indicate a gravitational anomaly in the low-acceleration regime, in line with the results of 2023-2024 based on $v_{p}$. This holds great promise for future studies with a significant number of precise values of $v_{r}$.
@@
This paper is structured as follows. In Section 2, I describe the methodology and data. In detail, I first develop a Bayesian methodology to infer PDFs of gravity from individual binaries (Section 2.1), and then describe the method of statistically combining individual PDFs to obtain a consolidated probability distribution by maximizing the information (Section 2.2). Section 2.3 describes the selection of a pure binary sample from the Gaia DR3 database that will be used for individual Bayesian modeling. The selection of the sample will be more stringent than previous samples that were used mainly for median statistics in bins of accelerations (or equivalent variables) by Chae (2024a, c). In Section 3, I describe the results of Bayesian modeling based on the Gaia sample. In Section 3.1, I first examine the individual Bayesian results and the overall properties of the results. In Section 3.2, I present the consolidated probability distributions of gravity with wide binaries in a Newtonian (relatively strong internal acceleration) regime. In Section 3.3, I present the main results of this work, i.e., the consolidated probability distributions of gravity in low-acceleration regimes. In Section 4, I discuss the meaning of the results and future wide binary gravity research. In Section 5, I present the current conclusion and prospects with future data on radial velocities (RVs). In Appendix A, a general transformation with the Euler angles between the orbital plane and the observer’s frame is described as it is needed to produce mock samples. In Appendix B, perspective effects in PMs and RVs of wide binary stars are investigated to explore any possible biases in the low-acceleration regime data.
@@ 2.1 Bayesian 3D modeling
A gravitational anomaly with respect to the prediction of Newtonian gravity can be defined and tested/measured in various ways. Because of the extremely long periods of wide binaries compared with the observational time span, only one relative velocity is observed in effect. In Newtonian gravitational dynamics, the relative motion between the pair in an isolated binary system follows an ellipse. In nonstandard theories of gravity such as MOND gravity (e.g., Bekenstein & Milgrom 1984; Milgrom 2010), the orbits of wide binaries under the external field of the Milky Way oriented at an arbitrary angle with respect to the orbital plane may not be exact ellipses. However, whether the actual orbits are exact ellipses or not, I will work with elliptical orbits. The assumption of elliptical orbits will have no impact on testing Newtonian gravity, i.e., whether a gravitational anomaly is present or not. However, when a gravitational anomaly is present, the quantified gravitational anomaly must be understood as an effective quantity for elliptical orbits.
@@
Consider Figure 1, which connects the observer’s primed frame with the unknown orbital plane. Note that the $x$ and $x^{\prime}$ axes coincide by the definition of the two coordinate systems in this figure. (See Appendix A for fully general transformations.) As in Gaia observations, suppose that observed quantities for the two stars of each binary are the R.A. ($\alpha_{k}$ with $k=A,B$ referring to the brighter and fainter components hereafter) and decl. ($\delta_{k}$) given in units of degree, PMs ($\mu^{\star}_{\alpha,k}$ and $\mu_{\delta,k}$) given in units of mas yr-1, and RVs ($\text{RV}_{k}$) given in units of km s-1. What matter for internal dynamics are the relative positions and relative velocities between the pair.
@@
which will be used as an observational constraint on the orbit. The relative RV is given by
@@
The above equations (Equations (5) and (7)) for the relative velocity components based on the observed PMs and RVs do not include corrections for the “perspective effects” (Shaya & Olling, 2011), which are apparent relative velocities between the pair in the observer’s spherical coordinates due to the motion of the barycenter of the system in the 3D space. For wide binaries in the solar neighborhood, perspective effects can be relevant for $s>0.1$ pc (i.e., 20 kau; El-Badry 2019).
Almost all binaries considered in this work have $s<20$ kau (99% satisfy $s<15.5$ kau) meaning that perspective effects are minimal, and thus their corrections are not warranted when RV uncertainties are large. Moreover, for an accurate estimate of the perspective effects, the difference in the distances of the two stars is needed. Unfortunately, observed distance uncertainties are (usually much) larger than $s$ and one has to assume zero (or reasonably small) difference in the distances to obtain a first-order estimate of the perspective effects. The default choice of the present work will be to ignore any perspective effects. However, I will also consider the case of correcting for first-order perspective effects estimated by a numerical procedure described in Appendix B.
I define the ratio of the RV to a projected velocity component as
@@
which is related to the “inclination” angle $i$ by $\tau_{\rm{obs}}=\tan i$ [^1] Here the parameter $i$ is the same as the angle between the orbital plane and the plane of the sky only in the special case of Figure 1. In general cases where the sky plane is rotated by an arbitary angle about the $z^{\prime}$-axis as shown Appendix A, $i$ is not the true inclination angle but a mere parameter. Finally, the magnitude of the relative physical velocity in the 3D space is then given by
@@
In Equations (10), (12), and (13), $e$ is the eccentricity, $i$ is the inclination parameter, $\phi_{0}$ is the longitude of the periastron, and $\phi$ is the phase angle given by
@@
where $\{X_{j,\rm{mod}}(\mathbf{\Theta})\}$ ($j=1,2,3$) are the model predictions given by Equations (10), (12), and (13) and $\{X_{j,\rm{obs}}\}$ are the three observational constraints given by Equations (9), (6), and (8). [^2] In principle, $\Delta d(=d_{B}-d_{A})$ can provide another observational constraint, but its observational uncertainty is usually much larger than the line-of-sight separation between the stars, so that it is not useful. Nevertheless, the Bayesian modeling code (Chae, 2025) has the option to include the constraint. In Equation (15), $\{\sigma_{j}\}$ are the observational uncertainties estimated as follows.
@@
where $v_{x^{\prime}}$ and $v_{y^{\prime}}$ are the relative projected velocity components given by Equation (4) while $\sigma_{v_{x^{\prime}}}$ and $\sigma_{v_{y^{\prime}}}$ are their respective uncertainties arising from PM uncertainties, and $v_{r}$ is the relative RV given by Equation (7) while $\sigma_{v_{r}}$ is its uncertainty arising from the uncertainties of RVA and RVB. The above uncertainties are based on the standard error propagation procedure.
@@
For eccentricity $e$ I consider the range $0.9990.001<e<$ 0.999 with a prior probability distribution. For the prior, I consider three choices: a flat prior, a “thermal” probability density $p(e)=2e$, or a separation($s$)-dependent probability density $p(e)=(1+\alpha)e^{\alpha}$ with $\alpha=\alpha(s)$ from Hwang et al. (2022) (hereafter referred to as the “H2022” prior) as described by Equation (18) of Chae (2024a). If the data can determine the orbit well, the flat prior may be sufficient. However, with snapshot data, $e$ is ill constrained in particular for relatively large uncertainties of $v_{r}$ as in many cases based on the current Gaia DR3 database. When $e$ is ill constrained (or essentially unconstrained in some cases), the prior matters because the flat prior would be equivalent to assuming that the median eccentricity is $0.5$ while the thermal prior has a median of $\sqrt{0.5}(\approx 0.71)$ more consistent with the observed statistical properties. The H2022 prior has $\alpha\approx 1.3$ for $s>1$ kau(kilo astronomical unit) implying a median of $\approx 0.74$ that differs from the thermal case by just $0.03$, indicating that two priors would make little difference. Moreover, because individual data themselves have some constraining power in this work, an eccentricity probability distribution plays a weaker role as a prior than a fixed input as in previous studies. For this reason, the thermal prior will be the default choice.
@@
For $f_{M}$, I assume a Gaussian probability density for $\log_{10}(f_{M})$ with $(\mu,\sigma)=(0,0.021)$ allowing a 5% scatter of $M_{\rm{tot}}$, which covers uncertainties of the mass estimates (see Section 2.3). For $i$ and $\Gamma$, I assume flat priors in the ranges: $-90^{\circ}<i<90^{\circ}$ and $-1<\Gamma<1$.
@@
In this subsection, the Bayesian method described in Section 2.1 is applied to mock Newtonian binary systems to investigate the properties of individual PDFs of $\Gamma$ (Equation (1)). Also, I investigate how individual PDFs of $\Gamma$ based on independent data (i.e., snapshot motions) can be combined to give a consolidated probability distribution.
For the orbital motion shown in Figure 1, consider binary systems with total mass $M_{\rm{tot}}=2.0M_{\odot}$, eccentricity $e=0.75$, and $\phi_{0}=-135^{\circ}$. For the semi-major axis of the elliptical orbit, I consider two cases: $a=0.8$ kau and $8$ kau to represent binaries in the Newtonian ($g_{\rm{N}}\gtrsim 10^{-8}$ m s-2) and MOND ($g_{\rm{N}}\lesssim 10^{-9.5}$ m s-2) regimes. Consider a viewing angle $i=60^{\circ}$ in the geometry of Figure 1. For the sake of illustration, assume that both binaries obey Newtonian gravity regardless of their separation (i.e., $\gamma_{g}=1$ or $\Gamma=0$).
@@
$v(r)$ $=$ $\sqrt{\frac{G_{\rm N}M_{\rm{tot}}}{r}\left(2-\frac{r}{a}\right)}$ (23)
$=$ $941.9\sqrt{\frac{M_{\rm{tot}}/M_{\odot}}{r/\text{kau}}\left(2-\frac{r}{a}\right)}\text{ m s}^{-1},$
@@
where [^4] The presence/absence of the factor $\pi$ varies, but for Equation (24) it is irrelevant.
@@
Figure 2: Each row shows a simulated Newtonian orbit with eccentricity $e=0.75$ and semimajor axis $a=0.8$ or 8 kau. All positions and velocities represent relative displacements and relative velocities between the pair of stars with a total mass of $M_{\rm{tot}}=2M_{\odot}$. See Figure 1 for the geometry and the meanings of the other parameters. For each orbit the colored big dots represent 20 mock snapshot observations with uniformly spaced $t/P$ where $t$ is the time from the passage of the periastron (at $\phi=\phi_{0}$) and $P$ is the period. Because of long periods only snapshot observations are possible. The sky-projected velocities $v_{x^{\prime}}$ and $v_{y^{\prime}}$ would be observed as PMs in real observations while the line-of-sight (radial) velocity $v_{z^{\prime}}$ would be observed by spectroscopic observations.
@@
Figure 3: Each row shows three examples of the PDFs of the pseudo-Newtonian model parameters including $\Gamma$ (Equation (1)) derived for the mock Newtonian (i.e., with $\Gamma=0$) data shown in Figure 2. For the Bayesian modeling, an uncertainty of 200 m s-1 is assumed for each radial velocity of a star, thus $\approx 283$ m s-1 for $v_{r}$.
@@
Figure 3 shows examples of Bayesian inference with the above procedure for the mock Newtonian data of the two binaries with $a=0.8$ kau or $8$ kau shown in Figure 2. These results represent some expectations for the current nominal Gaia statistical sample if gravity is Newtonian (or not too different from it). In general, the PDFs of the parameters except for $f_{M}$ are quite broad as expected considering the input sizable uncertainties (i.e., $\sigma_{v_{r}}\approx 283$ m s-1) of the relative RV ($v_{r}$). This is particularly true for the data of the binary with $a=8$ kau because its $v_{r}$ values have absolute values even smaller than $\sigma_{v_{r}}$ in some mock observation phases. The PDF of $\log_{10}(f_{M})$ is always equivalent to the prior Gaussian distribution because of its degeneracy with $\Gamma$, which has a flat prior in a broad range.
Although the individual PDFs of $\Gamma$ and the orbital and geometric parameters $e$, $\phi_{0}$, and $i$ are very broad, they are all consistent with the input values within reasonable confidence limits (e.g., $2\sigma$). Each column of Figure 4 shows the individual PDFs of $\Gamma$ for the mock data set of the Newtonian binary with $a=0.8$ kau or $8$ kau. The lower rows use the logarithmic scale for probability density to enhance visibility.